# Stabilizing and Solving Inverse Problems using Data and Machine Learning

### v1

@@ 6 Conclusions
The regularization of severely ill-posed inverse problems using large data sets and stabilized finite element methods was considered and shown to be feasible both for linear and non-linear problems. In the linear case, a fairly complete theory for the approach exists, and herein, we complemented previous work with the design and analysis of a reduced-order model. In the linear case, a combination of POD for the data reduction and reduced model method for the PDE-solution was shown to be a rigorous and robust approach that effectively can improve stability from logarithmic to linear in the case where the data is drawn from some finite dimensional space of moderate dimension. To extend the ideas to non-linear problems we introduced a machine learning framework, both for the data compression and the reduced model. After successful training, this resulted in a very efficient method for the solution of the non-linear inverse problem. The main observations were the following:
@@
    Machine learning can be used to leverage the observations in the linear case to non-linear inverse problems and data assimilation and results in fast and stable reconstruction methods.
@@ Acknowledgements.
This research was supported in part by the Swedish Research Council Grants No.  2021-04925, and the Swedish Research Programme Essence. EB acknowledges funding from EP/T033126/1 and EP/V050400/1.

### v2

@@ 6 Conclusions
The regularization of severely ill-posed inverse problems using large data sets and stabilized finite element methods was considered and shown to be feasible both for linear and nonlinear problems. In the linear case, a fairly complete theory for the approach exists, and herein, we complemented previous work with the design and analysis of a reduced-order model. In the linear case, a combination of POD for the data reduction and reduced model method for the PDE-solution was shown to be a rigorous and robust approach that effectively can improve stability from logarithmic to linear in the case where the data is drawn from some finite-dimensional space of moderate dimension. To extend the ideas to nonlinear problems we introduced a machine learning framework, both for the data compression and the reduced model. After successful training, this resulted in a very efficient method for the solution of the nonlinear inverse problem. The main observations were the following:
@@
    Machine learning can be used to leverage the observations in the linear case to nonlinear inverse problems and data assimilation and results in fast and stable reconstruction methods.
@@ Acknowledgements.
This research was supported in part by the Swedish Research Council Grant, No.  2021-04925, and the Swedish Research Programme Essence. EB acknowledges funding from EPSRC grants EP/T033126/1 and EP/V050400/1.